# IncogniText: Privacy-enhancing Conditional Text Anonymization via LLM-based Private Attribute Randomization

### v1

@@ 4 Conclusion
5 Limitations
While our method achieves tremendous reduction of the private attribute attacker accuracy, the attacker might use a stronger attribute inference model, e.g., a model finetuned for this task, than the open-source adversary model we used in our experiments. This is especially true for on-device setting as the adversary model used has to also be on-device and therefore must be small, e.g., Qwen 1.5B in our experiments. Using better models, e.g., GPT-4, for privacy evaluation might also reveal a higher privacy leakage. Nevertheless, we believe IncogniText would still achieve substantially higher protection than the baselines. Finally, conducting the utility evaluation with humans, e.g., with Likert score Likert (1932), would yield more insightful results into the willingness of people to use this technique in a real-world application.
@@ Appendix A Ablation results
As mentioned in Section 3, we present the following ablation results on models other than Phi-3-small as anonymizers.
@@
Table 4: Results (%) of the ablation study conducted with Llama-3-70B as anonymization model and evaluated with Phi-3-small.
@@
Table 5: Results (%) of the ablation study conducted with Llama-3-8B as anonymization model and evaluated with Phi-3-small.
@@
Table 6: Results (%) of the ablation study conducted with Phi-3-mini as anonymization model and evaluated with Phi-3-small.
Appendix B Preprocessing of the self-disclosure dataset
As mentioned in Section 3, we use the self disclosure dataset from Dou et al. (2023) as a starting point. We consider the following attributes: gender, relationship status, age, education, and occupation. We keep only samples where the author discloses information about their own private attributes and not about someone else. Furthermore, we label the samples with the real private attribute values instead of text spans, yielding a set of 196 examples.
Appendix C Finetuning details
We provide further details to the finetuning data and process. First, we construct the finetuning dataset based on samples from the synthetic conversations in Staab et al. (2023) that were not included in the officially released set. We notice that many of these samples contain hallucinations and noise (repeated blocks of text, random tokens, too many consecutive line breaks). We filter these samples out. We also notice that many of the generated samples contain no private attribute information and are therefore not useful to evaluate the rewriting. Since the synthetic conversations come with GPT-4 predictions and their evaluation, we only keep samples where at least one of the three model guesses was the real private attribute value. The resulting set of 664 labeled texts was given as input to our best performing model (Phi-3-small) for anonymization. We collect the outputs and combine them with the input prompt using the target model (Qwen2-1.5B) template. The resulting dataset is the one we use for instruction finetuning. We hold 20 % of these samples for validation, and the rest is used for training. We use bi-gram ROUGE for evaluation.
@@ Appendix C Finetuning details
Appendix D Additional results
We present further results showcasing the differences between anonymization with and without a target attribute value. Figure 2 is a histogram showing that more than 80% of samples are already anonymized in the first iteration using our method, wheras more than half of samples need to go through a second and possibly a third iteration in FgAA.
Appendix E Prompt templates
The following are the prompt templates used for the anonymizer (conditioned on inference, ground truth and target value) and for the adversary. Similar to Staab et al. (2023), we use a format correction prompt to avoid parsing failures when the model doesn’t give the answer in the expected format. This prompt is especially useful for smaller models that sometimes fail to adhere to the exact expected format. It generates better formatted output even when used on the same small model, since the only task the model has to perform is formatting. We also use Staab et al. (2023)’s model aided evaluation prompt to decide whether the prediciton of the anonymizer is correct, for attributes where exact string matching is too restrictive (Example: ’Bachelors in Computer Science’ and ’B.Sc Computer Science’). We also include the LLM-based utility judge template used in Staab et al. (2024).

### v2

@@ Appendix A Ablation results
As mentioned in Section 3, we present the following ablation results on models other than Phi-3-small [2] as anonymizers.
@@
Table 4: Results (%) of the ablation study conducted with Llama-3-70B [16] as anonymization model and evaluated with Phi-3-small [2].
@@
Table 5: Results (%) of the ablation study conducted with Llama-3-8B [16] as anonymization model and evaluated with Phi-3-small [2].
@@
Table 6: Results (%) of the ablation study conducted with Phi-3-mini [16] as anonymization model and evaluated with Phi-3-small [2].
Appendix B Finetuning details
We provide further details to the finetuning data and process. First, we construct the finetuning dataset based on samples from the synthetic conversations in [22] that were not included in the officially released set. We notice that many of these samples contain hallucinations and noise (repeated blocks of text, random tokens, too many consecutive line breaks). We filter these samples out. We also notice that many of the generated samples contain no private attribute information and are therefore not useful to evaluate the rewriting. Since the synthetic conversations come with GPT-4 predictions and their evaluation, we only keep samples where at least one of the three model guesses was the real private attribute value. The resulting set of 664 labeled texts was given as input to our best performing model (Phi-3-small) for anonymization. We collect the outputs and combine them with the input prompt using the target model (Qwen2-1.5B) template. The resulting dataset is the one we use for instruction finetuning. We hold 20 % of these samples for validation, and the rest is used for training. We use bi-gram ROUGE for evaluation.
@@ Appendix B Finetuning details
Appendix C Prompt templates
The following are the prompt templates used for the anonymization model (conditioned on inference, ground truth and target value) and for the adversary. Similar to [22], we use a format correction prompt to avoid parsing failures when the model doesn’t give the answer in the expected format. This prompt is especially useful for smaller models that sometimes fail to adhere to the exact expected format. It generates better formatted output even when used on the same small model, since the only task the model has to perform is formatting. We also use [22]’s model aided evaluation prompt to decide whether the prediciton of the anonymization model is correct, for attributes where exact string matching is too restrictive (Example: ’Bachelors in Computer Science’ and ’B.Sc Computer Science’). We also include the LLM-based utility judge template used in [23].